# Enhancing Screen Time Identification in Children with a Multi-View Vision Language Model and Screen Time Tracker

### v2

@@ 1 Introduction
Figure 1: The STT device (left panel); Compilation of images showcasing various environments in screen time exposure. The montage is created using free-living data collected from STT (right panel). The STT device is lightweight and can be firmly attached to clothes.
@@
A wearable camera with the capability to capture egocentric images is a promising non-invasive candidate to monitor screen presence. Egocentric images are captured from first-person perspective and can demonstrate the wearer’s viewpoint. Although there is existing work that uses wearable cameras to track children’s screen exposure [10, 11], those solutions are not ideal due to their bulky size and lack of automated analysis. Artificial intelligence (AI) enables computers to automatically identify objects with high accuracy, making it promising for screen exposure measurement. While AI can identify object types from egocentric images, no automated algorithm has been customized for children’s screen exposure. Recently, vision language models have shown great potential for robustly identifying objects and events from images [12], but no such efforts have been made specifically for egocentric images for screen type identification. Convolutional neural network (CNN), as a major branch of AI, has advanced automated object detection methods in various applications by providing the backbone architectures for the development of the regional based CNN (R-CNN)[13] and You Only Look Once (YOLO)[14] detection systems. However, such methods work on individual frames and relies on local features extracted from CNN, lacking the capability to associate long range dependency of features or high-level objects.
@@
In this paper, we proposed a combination of an egocentric camera, namely the screen time tracker (STT), and a vision language model (VLM) to identify screen existence among multiple screen devices. The STT device and representative images are shown in Fig. 1. In particular, we devised a multi-view VLM (MV-VLM) to process sequential frames from egocentric images. We took advantage of a customized wearable sensor and egocentric images to identify children’s screen exposure. In comparison with existing single view input vision language models, we developed a unique multi-view vision language model that process egocentric image sequences from wearable devices. Notably, a contrastive learning-based view selection module and a screen type identification module are specifically designed to address the challenges in existing method and to develop a robust approach to detecting children’s screen exposure. The proposed approach innovatively takes multi-view rather than single-view images as input compared to the existing screen identification work [15, 16]. The innovation contributes to the model’s capability in extracting long-dependent textual features alongside spatial-temporal features from multi-view imaging, achieving the highest performance among all three types of screen scenarios.
@@
  • We proposed a wearable sensor solution to identifying children’s screen exposure. Our lightweight wearable sensor is children-friendly and captures egocentric images to monitor children’s electronic screen activities. We created a dataset collected from children’s free-living activities over two days, the first of its kind for sensor informatics.
@@
  • We devised a MV-VLM model, which is specifically for analyzing image sequences from wearable cameras. To leverage temporal variations and capture more spatial features for accurate type identification, we developed a novel model that takes multi-view inputs and fuses features from different views for language generation. The selection of multi-view images is based on contrastive learning, which maximizes the information among images from different views.
@@ 2 Related Work
Multi-view based Neutral Networks. A multi-view deep learning model leverages data from multiple perspectives to enhance learning in tasks such as classification and recognition by integrating diverse sources of information. Although it has been demonstrated in exsiting studies that the incorporation of multi-view in CNN-based models outperforms models with single-view [24, 25], multi-view network has not been fully explored in processing egocentric image sequences. A critical problem in applying multi-view network is how to select multi-view images from egocentric image sequences.
Large language model (LLM) in Egocentric Videos. Large language model is a rising field and incorporating LLMs into image processing has the potential to significantly improve upon the studies using traditional models such as CNNs. By leveraging the contextual capabilities of LLMs, researchers can achieve more accurate and comprehensive egocentric image understanding. LifelongMemory was proposed in [26] as a novel framework that uses multiple pre-trained models to answer queries from egocentric video content. Research in [27] addresses Ego4D natural language queries challenge with image and video captioning models. Most of the studies focuses on answer queries, thus lacking the capability to systematically analyze daily life for specific applications. Moreover, there is limited vision language model to address the identification of electronic screen type, even from a single view.
Figure 2: Proposed pipeline to process egocentric image sequences. During training, there are four major components: view selection, vision language model, language model, and screen identification. View selection module uses Contrastive Language-Image Pre-Training (CLIP) to extract embeddings and select multi-view images based on similarity. Vision Langauge model learns from vision transformer and MiniLM to generate textual description on multi-view images. In inference phase, only view selection, vision model, and screen identification model are used.
@@ 3.1 Wearable sensor and data collection
We collected data from an egocentric device, screen time tracker (STT), following our previous work [28]. The key component used in the sensor was a miniature 5 Megapixel camera with 120-degree wide-angle gaze-aligned lens with the resolution of $2592\times 1944$. The camera was attached to clothing with a magnetic clip mount, as shown in Fig. 1 (left panel), and boasted a battery life exceeding 48 hours. We redesigned the device’s casing by resembling a badge to reduce user burden, particularly for children. Positioned on the chest, this badge-like device captured egocentric images every 10 seconds. Our device measured a dimension of $54mm\times 35mm\times 12mm$. Notably, the device only takes 30% of the size of commercial cameras used in other wearable senor informatics work [10].
Young children between age of 3 and 5 were eligible to participate in data acquisition. The experimental protocol was approved by the local Institutional Review Board (IRB) at home institution. We collected data from 30 participants. Each participant wore the device for two days. The whole dataset roughly includes children’s participation in free-living activities. After the two day activities, a survey on the comfort of the wearable device is conducted following a conventional protocol [29]. The screen exposures related to TV, computers, and smartphones were captured. Representative images are shown in Fig. 1 (right panel). Image data were manually labeled by an undergraduate research assistant.
@@ 3.2 Framework
Our framework collected multiple images from different views to identify the existence of electronic screens. As shown in Fig. 2, we developed a novel screen identification framework that consists of a view selection module, a vision model, a language model, and a screen identification module. The core concept involves selecting and inputting images from multiple egocentric viewpoints to enhance the robustness of electronic screen identification.
@@ 3.2.1 Conceptual Rationale for Multi-View Vision Language Model
In the screen exposure scenarios, multi-view image processing could be beneficial in several aspects. For instance, when screens are only partially captured, multi-view images may provide complementary data from different parts of the screen, enabling deep learning models to better understand and integrate the screen’s various structural components. Additionally, in cases of low-quality images, such as those that are blurry, it is challenging for deep learning models to classify the images. Utilizing multi-view images allows the model to synthesize features from various perspectives, thereby enhancing feature representation and improving classification accuracy. To identify the existence of electronic screens, it is essential to consider the spatial relationships among static objects like TVs, computers, windows, walls, and shelves. These objects maintain consistent spatial relationships in the real world, which are reflected in 2D images based on perspective principles. To capitalize on these features and their spatial interrelations, we employed a vision transformer coupled with a language model. This approach leveraged the model’s capability to explore long-range dependencies in text descriptions and image processing.
@@ 3.2.2 View Selection
Image sequences captured from STT are in time series, representing objects from different views and timestamps. Images taken from the same scene with small variations of position, height, and orientation are considered similar, though environmental condition like light, time of the day may change. We sought for a mapping rule that is robust enough to reveal spatial relationships and partial or occluded screen information. To capture coherent image features, we chose a contrastive learning-based embedding approach to encode the images and analyze their similarity. Assuming the egocentric image sequences $I_{1},I_{2},I_{3},...,I_{n}$, for each image $I_{i}$, we used contrastive language-image pre-training (CLIP) to convert images to embeddings $CLIP(I_{i})$. For any two images $I_{m}$ and $I_{n}$, we measured the cosine similarity between two embeddings via:
@@
Where ($\cdot$) represents dot product and |||||$\cdot$||| | ⋅ | | is the magnitude. Then to split the image sequences into multi-view image groups, we built a graph $G(V,E)$ to represent the images with their similarity according to values of $Sim(I_{m},I_{n})$. In Graph $G$, each node $v_{i}$ represents an image $I_{i}$. $V$ corresponds to the set of nodes (i.e., images) and $E$ corresponds to the set of weighted edges that connect the each pair of nodes within a time window. For two nodes $v_{m}$ and $v_{n}$, the edge $E_{m,n}$ of them is $valid$ only if $Sim(I_{m},I_{n})$ is within a range between an upper bound $\tau_{h}$ and a lower bound $\tau_{l}$. The upper bound ensured that static images in consecutive frames would not be selected. The lower bound ensured these two images could capture the same screen-related scene. We built an undirected graph and selected connected components of size $k$ as multi-view image groups. To get the most number of multi-view groups, we first sorted the all connected components of size $k$ according to the sum of edge degrees in each $k$-size component ascendingly. Then we greedily selected and split the components starting from the components with least edge degrees from $G$, and updated $G$ accordingly. The detailed procedures are illustrated in Algorithm 1.
@@ 3.2.3 Vision Language Model
Text Embedding (MiniLM). In the training phase, we employed a MiniLM [32] for extracting textual features from annotations associated with multi-view images. In particular, each images were associated with a sentence of textual annotation during training. MiniLM is made of a set of distilled self-attention model in transformer implementation. In this study, our analysis was constrained by a limited sample space, as all egocentric images were sourced from indoor home environments. Furthermore, the sample set mainly comprises screen-related instances. In addition, textual information involved in this study was mainly that relevant to screen events. This made MiniLM particularly well-suited for our application, as its distilled architecture efficiently captured the necessary textual embeddings during the training phase. In the Feature Synthesis layer, text embeddings were normalized and concatenated to ensure the brevity of image captions. These concatenated embeddings were then processed through a linear alignment layer, ensuring dimensional consistency with the embeddings used in the Swin Transformer model. In particular, the text embeddings extracted from this phase serve as the annotation for the following text generation model.
@@
Figure 4: a) Text generation results from VLM to compare how close the generated scene description is similar to the ground truth annotation; b) Comparison among proposed method (MV-VLM), (YOLOv8) [34] and single-view vision language model (MiniGPT) [35] in overall accuracy; c) Comparison among proposed method (MV-VLM), (YOLOv8) [34] and single-view vision language model (MiniGPT) [35] in identification of TV, smartphone, and computer. Our methods consistently shows a higher performance than the other two methods.
@@ 4 Experiments and Results
Figure 5: Examples of generated text and screen identification. The left panel shows typical multi-view images. The right panel shows the generated description from language model, the screen identification results, and the ground truth. The generated scene description is logical and smooth, with key words that could be efficiently processed to categorize to specific screen type.
Table 2: Ablation study on the design of multi-view and multi-caption model in proposed design. We compared proposed method with implementation without the module of multi-view (MV) and/or CLIP-based view selection. The best scores are highlighted in red and the second best scores are highlighted in blue.
Method Multi-View (MV) CLIP Accuracy
TV Smartphone Computer
Proposed w/o MV, CLIP 0.77 0.84 0.75
Proposed w/o CLIP 0.81 0.79 0.74
Proposed 0.93 0.85 0.85
@@ 4.1 Dataset
The raw dataset has 1191 images, corresponding to 397 groups of multi-view images, including different screen types of TV, smartphone, computer use in free-living environments. To build a multi-view image dataset with caption describing the screen of each group, labels were manually generated with guidance from Bootstrapping Language-Image Pre-training (BLIP2) [36]. We divided the whole image set to four folds and conducted cross-validation to validate the performance of identifying various screen types.
@@ 4.2 Training strategy for implementation
The training process required minimal effort to customize the VLM for screen detection. In particular, we chose a pre-trained model in Swin Transformer for ViT [30] and MiniLM for language model [32]. We fine-tuned ViT using our multi-view image datasets. In text generation, Llama model ($Llama2-7B$ [33]) remained frozen as well. The training remained on the parameters in alignment layer. The experiments were conducted in parallel on three RTX A6000 GPUs. We set the batch size as 6 and a learning rate as 0.0001. In multi-view selection, the size $k$ of multi-view group was set as 3. We also empirically set $\tau_{h}$ as 70% and $\tau_{l}$ as 40%.

### v3

@@ 1 Introduction
Figure 1: a) The STT device (left panel); Compilation of images showcasing various environments in screen time exposure. The montage is created using free-living data collected from STT (right panel). The STT device is lightweight and can be firmly attached to clothes. b) The architecture diagram of the proposed MV-VLM. The MV-VLM is designed to process egocentric camera frames collected by a).
@@
A wearable camera with the capability to capture egocentric images is a promising non-invasive candidate to monitor screen presence. Egocentric images are captured from first-person perspective and can demonstrate the wearer’s viewpoint. Egocentric images have been utilized for human activity analysis [10, 11, 12]. Although there is existing work that uses wearable cameras to track children’s screen exposure [13, 14], those solutions are not ideal due to their bulky size and lack of automated analysis. Artificial intelligence (AI) enables computers to automatically identify objects with high accuracy, making it promising for screen exposure measurement. While AI can identify object types from egocentric images, no automated algorithm has been customized for children’s screen exposure. Recently, vision language models have shown great potential for robustly identifying objects and events from images [15, 16, 17]. However, no such efforts have been made specifically for egocentric images for screen type identification. Convolutional neural network (CNN), as a major branch of AI, has advanced automated object detection methods in various applications by providing the backbone architectures for the development of the regional based CNN (R-CNN)[18] and You Only Look Once (YOLO)[19] detection systems. However, such methods work on individual frames and relies on local features extracted from CNN, lacking the capability to associate long-range dependency of features or high-level objects.
@@
In this paper, we proposed a combination of an egocentric camera, namely the screen time tracker (STT), and a vision language model (VLM) to identify screen existence among multiple screen devices. The STT device and representative images are shown in Fig. 1 a). In particular, we devised a multi-view VLM (MV-VLM) to process sequential frames from egocentric images as shown in Fig. 1 b). Large language models provide a human-readable explanation to the insights of images compared with traditional classification methods. In addition, it unveils the spatial relationship among multiple objects that are relevant to the existence of the screen. Particularly, with the involvement of transformers in LLM, long-range dependence of local features is taken into consideration. Moreover, the multi-view scheme also explores the temporal features, making our proposed model particularly capable of detecting a screen that is in occlusion. We took advantage of a customized wearable sensor and egocentric images to identify children’s screen exposure. In comparison with existing single view input vision language models, we developed a unique multi-view vision language model that process egocentric image stream from wearable devices. Notably, a contrastive learning-based view selection module and a screen type identification module are specifically designed to address the challenges in existing method and to develop a robust approach to detecting children’s screen exposure. The proposed approach innovatively takes multi-view rather than single-view images as input compared to the existing screen identification work [20, 21]. The innovation contributes to the model’s capability in extracting long-dependent textual features alongside spatial-temporal features from multi-view imaging, achieving the highest performance among all three types of screen scenarios.
@@
  • We proposed a wearable sensor solution to identifying children’s screen exposure. Our lightweight wearable sensor is children-friendly and captures egocentric images to monitor children’s electronic screen activities. We created a dataset collected from children’s free-living activities over two days, the first of its kind for egocentric camera.
@@
  • We devised a MV-VLM model, which is specifically for analyzing image stream from wearable cameras. To leverage temporal variations and capture more spatial features for accurate type identification, we developed a novel model that takes multi-view inputs and fuses features from different views for language generation. The selection of multi-view images is based on contrastive learning, which maximizes the information among images from different views.
@@ 2 Related Work
Multi-view based Neutral Networks. A multi-view deep learning model leverages data from multiple perspectives to enhance learning in tasks such as classification and recognition by integrating diverse sources of information. Its potential has been explored in recent research [29, 30]. Although it has been demonstrated in existing studies that the incorporation of multi-view in CNN-based models outperforms models with single-view [31, 32], multi-view network has not been fully explored in processing egocentric image stream. A critical problem in applying multi-view network is how to select multi-view images from egocentric image stream.
Large language model (LLM) in Egocentric Streams. Large language model is a rising field and incorporating LLMs into image processing has the potential to significantly improve upon the studies using traditional models such as CNNs. By leveraging the contextual capabilities of LLMs, researchers can achieve more accurate and comprehensive egocentric image understanding. LifelongMemory was proposed in [33] as a novel framework that uses multiple pre-trained models to answer queries from egocentric video content. Research in [34] addresses Ego4D natural language queries challenge with image and video captioning models. Most of the studies focuses on answer queries, thus lacking the capability to systematically analyze daily life for specific applications. Moreover, there is limited vision language model to address the identification of electronic screen type, even from a single view.
Figure 2: Proposed pipeline to process egocentric image stream. There are four major components: view selection, vision language model, language model, and screen identification. View selection module uses Contrastive Language-Image Pre-Training (CLIP) to extract embeddings and select multi-view i mages based on similarity. Vision Langauge model learns from vision transformer and MiniLM to generate textual description on multi-view images.
@@ 3.1 Wearable sensor and data collection
We collected data from an egocentric device, screen time tracker (STT), following our previous work [35]. The key component used in the sensor was a miniature 5 Megapixel camera with 120-degree wide-angle gaze-aligned lens with the resolution of $2592\times 1944$. The camera was attached to clothing with a magnetic clip mount, as shown in Fig. 1 a) (left panel), and boasted a battery life exceeding 48 hours. We redesigned the device’s casing by resembling a badge to reduce user burden, particularly for children. Positioned on the chest, this badge-like device captured egocentric images every 10 seconds. Our device measured a dimension of $54mm\times 35mm\times 12mm$. Notably, the device only takes 30% of the size of commercial cameras used in other wearable senor informatics work [13].
Young children between age of 3 and 5 were eligible to participate in data acquisition. The experimental protocol was approved by the local Institutional Review Board (IRB) at home institution. We collected data from 30 participants. Each participant wore the device for two days. The whole dataset roughly includes children’s participation in free-living activities. After the two day activities, a survey on the comfort of the wearable device is conducted following a conventional protocol [21, 20]. The screen exposures related to TV, computers, and smartphones were captured. Representative images are shown in Fig. 1 a) (right panel). Image data were manually labeled by an undergraduate research assistant.
@@ 3.2 Framework
Our framework collected multiple images from different views to identify the existence of electronic screens. As shown in Fig. 2, we developed a novel screen identification framework that consists of a view selection module, a vision model, a language model, and a screen identification module. The core concept involves selecting and inputting images from multiple egocentric viewpoints to enhance the robustness of electronic screen identification. The language model unveils the spatial relationship among multiple objects that are relevant to the existence of the screen. Particularly, with the involvement of transformers in LLM, long-range dependence of local features is highlighted. The Multi-View scheme also explores the temporal features, making our proposed model particularly capable of detection screen that is in occlusion.
@@ 3.2.1 Conceptual Rationale for Multi-View Vision Language Model
As shown in Fig. 2, due to the motion of children during an screen event, the egocentric image stream usually includes views at various angles to the electronic screen. In the screen exposure scenarios, Multi-View image processing could be beneficial in several aspects. For instance, when screens are only partially captured, Multi-View images may provide complementary data from different parts of the screen, enabling deep learning models to better understand and integrate the screen’s various structural components. Additionally, in cases of low-quality images, such as those that are blurry, it is challenging for deep learning models to classify the images. Utilizing Multi-View images allows the model to synthesize features from various perspectives, thereby enhancing feature representation and improving classification accuracy. To identify the existence of electronic screens, it is essential to consider the spatial relationships among static objects like TVs, computers, windows, walls, and shelves. These objects maintain consistent spatial relationships in the real world, which are reflected in 2D images based on perspective principles. To capitalize on these features and their spatial interrelations, we employed a vision transformer coupled with a language model. This approach leveraged the model’s capability to explore long-range dependencies in text descriptions and image processing.
@@ 3.2.2 View Selection
The image stream captured from STT is in time series, representing objects from different views and timestamps. The term ’view’ used in this paper is similar to ’perspective’. It refers to images captured from the same camera at different time instants, but within the same scene. This view selection module converts temporal change in images frames to spatial change in perspective/view. Images taken from the same scene with small variations of position, height, and orientation are considered similar, though environmental condition like light, time of the day may change. We sought for a mapping rule that is robust enough to reveal spatial relationships and partial or occluded screen information. To capture coherent image features, we chose a contrastive learning-based embedding approach to encode the images and analyze their similarity. Assuming the egocentric image stream $I_{1},I_{2},I_{3},...,I_{n}$, for each image $I_{i}$, we used contrastive language-image pre-training (CLIP) [36] to convert images to embeddings $CLIP(I_{i})$. The CLIP enables pairwise image similarity comparison without prior knowledge, making it more adaptable than heuristic methods such as K-Means, which rely on predefined clusters and struggle with the variability of real-world screen detection images.
For any two images $I_{m}$ and $I_{n}$, we measured the cosine similarity between two embeddings via:
@@
Where ($\cdot$) represents dot product and |||||$\cdot$||| | ⋅ | | is the magnitude. Then to split the image stream into Multi-View image groups, we built a graph $G(V,E)$ to represent the images with their similarity according to values of $Sim(I_{m},I_{n})$. In Graph $G$, each node $v_{i}$ represents an image $I_{i}$. $V$ corresponds to the set of nodes (i.e., images) and $E$ corresponds to the set of weighted edges that connect the each pair of nodes within a time window. For two nodes $v_{m}$ and $v_{n}$, the edge $E_{m,n}$ of them is $valid$ only if $Sim(I_{m},I_{n})$ is within a range between an upper bound $\tau_{h}$ and a lower bound $\tau_{l}$. The upper bound ensured that static images in consecutive frames would not be selected. The lower bound ensured these two images could capture the same screen-related scene. We built an undirected graph and selected connected components of size $k$ as Multi-View image groups. To get the most number of Multi-View groups, we first sorted the all connected components of size $k$ according to the sum of edge degrees in each $k$-size component ascendingly. Then we greedily selected and split the components starting from the components with least edge degrees from $G$, and updated $G$ accordingly. The detailed procedures are illustrated in Algorithm 1.
@@ 3.2.3 Vision Language Model
Text Embedding (MiniLM). In the training phase, we employed a MiniLM [39] for extracting textual features from annotations associated with Multi-View images. In particular, each images were associated with a sentence of textual annotation during training. MiniLM is made of a set of distilled self-attention model in transformer implementation. In this study, our analysis was constrained by a limited sample space, as all egocentric images were sourced from indoor home environments. Furthermore, the sample set mainly comprises screen-related instances. In addition, textual information involved in this study was mainly that relevant to screen events. This made MiniLM particularly well-suited for our application, as its distilled architecture efficiently captured the necessary textual embeddings during the training phase. In the Feature Synthesis layer, text embeddings were normalized and concatenated to ensure the brevity of image captions. These concatenated embeddings were then processed through a linear alignment layer, ensuring dimensional consistency with the embeddings used in the Swin Transformer model. In particular, the text embeddings extracted from this phase serve as the annotation for the following text generation model. We adopt the MiniLM for processing the text input for two reasons: MiniLM maps input texts of varying lengths into fixed-length embeddings, enhancing the robustness and scalability of the proposed framework. Moreover, by using the embeddings as input, the proposed framework can focus on high-level semantic information from the text.
Figure 3: Details of dataset acquired in this research. a) The overall distribution of different screen types. b) The number of image groups acquired from each subject.
Alignment layer. The alignment layer is a fully-connected layer which projects embeddings of multi-modal data to the LLMs’ input space. The fused multi-modal data by concatenation is sent to the alignment layer. Then processed embeddings are sent to the LLM for screen identification. The alignment layer has trainable parameters that can be updated during the training process, in order to achieve the optimal projection of the multi-modal data to the LLM’s input space.
@@ 4.1 Dataset
The dataset has 1,800 images, corresponding to 600 groups of Multi-View images, including different screen types of TV, smartphone, computer use in free-living environments. The condition of the free-living environment includes scenarios where light is bright or dark. The detailed distribution of the screen types is shown in Fig. 3, which aligns with the screen usage distribution in the study [24]. Moreover, we provide the number of image group acquired from each subject. To build a Multi-View image dataset with caption describing the screen of each group, labels were manually generated with guidance from Bootstrapping Language-Image Pre-training (BLIP2) [41] and Indoor Scene Understanding [42]. The captions will provide environment information besides screens. The spatial relationship between screens and the environments will be provided and extracted by the language model for screen detection. We divided the whole image set to four folds based on random shuffle over candidates and conducted cross-validation to validate the performance of identifying various screen types.
Figure 5: Examples of generated text and screen identification. The left panel shows typical Multi-View images. The right panel shows the generated description from language model, the screen identification results, and the annotation. Key words could be efficiently processed to categorize to specific screen type.
@@ 4.2 Training strategy for implementation
The training process required minimal effort to customize the VLM for screen detection. In particular, we chose a pre-trained model in Swin Transformer for ViT [37] and MiniLM for language model [39]. We fine-tuned ViT using our Multi-View image datasets. In text generation, Llama model ($Llama2-7B$ [40]) remained frozen as well. The training remained on the parameters in alignment layer. The experiments were conducted in parallel on two NVIDIA H100 GPUs. We set the batch size as 6 and a learning rate as 0.0001. In Multi-View selection, the size $k$ of Multi-View group was set as 3. We also empirically set $\tau_{h}$ as 70% and $\tau_{l}$ as 40%.
@@ 4.3 Multi-View selection
Figure 6: a) Text generation results from proposed method (MV-VLM) and LLaVA to compare how close the generated scene description is similar to the ground truth annotation; b) Comparison among MV-VLM, CLIP-fC, LLaVA, YOLOv11, and MV-Swin-T in overall accuracy; c) Comparison among MV-VLM, CLIP-fC, LLaVA, YOLOv11, and MV-Swin-T in identification of TV, smartphone, and computer. Our methods consistently shows a higher performance than the other baseline methods.